OpenStudy (anonymous):

If z=14+y, for y and z=10 then y=? what is the expression too?

OpenStudy (anonymous):

If I understand the question correctly, you have z=14+y and therefore y=z-14. If z=10, then plugging that into the equation, you get y=-4.